### v3

# CoVR: Learning Composed Video Retrieval from Web Video Captions

Abstract
Composed Image Retrieval (CoIR) has recently gained popularity as a task that considers both text and image queries together, to search for relevant images in a database. Most CoIR approaches require manually annotated datasets, comprising image-text-image triplets, where the text describes a modification from the query image to the target image. However, manual curation of CoIR triplets is expensive and prevents scalability. In this work, we instead propose a scalable automatic dataset creation methodology that generates triplets given video-caption pairs, while also expanding the scope of the task to include composed video retrieval (CoVR). To this end, we mine paired videos with a similar caption from a large database, and leverage a large language model to generate the corresponding modification text. Applying this methodology to the extensive WebVid2M collection, we automatically construct our WebVid-CoVR dataset, resulting in 1.6 million triplets. Moreover, we introduce a new benchmark for CoVR with a manually annotated evaluation set, along with baseline results. Our experiments further demonstrate that training a CoVR model on our dataset effectively transfers to CoIR, leading to improved state-of-the-art performance in the zero-shot and finetuning setups on both the CIRR and FashionIQ benchmarks. Our code, datasets, and models are publicly available at https://imagine.enpc.fr/~ventural/covr.
1 Introduction
Consider the scenario where a traveller takes a picture of a landmark or scenic spot and wants to discover videos that capture the essence of that location, by specifying certain conditions via text. For example, the query image in Figure 1 (of a fountain in Barcelona), along with the text “during show” should bring the video showcasing the fountain show. Further refining the text query such as “during show at night”, would allow the traveller to decide whether to wait for the show until the night time. In this work, our goal is composed video retrieval (CoVR), where the user performs such multi-modal search, by querying an image of a particular visual concept and a modification text, to find videos that exhibit the similar visual characteristics with the desired modification, in a dynamic context.
Figure 1: Task: Composed Video Retrieval (CoVR) seeks to retrieve videos from a database by searching with both a query image and a query text. The text typically specifies the desired modification to the query image. In this example, a traveller might wonder how the photographed place looks like during a fountain show, by describing several modifications, such as “during show at night, with fireworks”.
CoVR has many use cases, including but not limited to searching online videos for finding reviews of a specific product, how-to videos of a tool for specific usages, live events in specific locations, sports matches of specific players. Similar to composed image retrieval (CoIR), CoVR is also particularly useful when conveying a concept with a visual is easier and/or more accurate than only using words (e.g., unknown location/object, specific camera view, specific color).
Given the increased momentum in vision and language research in the recent years (Li et al. 2022b; Radford et al. 2021), CoIR has emerged as a new task (Vo et al. 2019), and since then witnessed improvements of both models and benchmarks (Baldrati et al. 2023, 2022; Gu et al. 2023; Levy et al. 2024; Liu et al. 2021; Wu et al. 2021). However, to the best of our knowledge, CoVR was not studied before. A key challenge in building CoVR models is the difficulty of gathering suitable training data of video-text-video triplets. We overcome this limitation by developing an automatic approach to generate triplets from existing video-caption collections. Specifically, we mine video pairs whose corresponding captions slightly differ in text space. We automatically describe this difference with a language model, which we train for a modification-text generation task. In particular, we use manually annotated triplets, each containing: (a) source caption, (b) target caption, (c) the modification text. We then finetune a large language model (LLM) (Touvron et al. 2023) by inputting (a-b), and outputting (c). We assume the resulting modification to describe the difference between the corresponding videos, thus obtaining video-text-video triplets (see Figure 2 for an overview). When training our CoVR/CoIR models, we can flexibly select one or more frames from the videos, enabling multiple settings (i.e., retrieving images or videos).
Figure 2: Method overview: We automatically mine similar caption pairs from a large video-caption database from the Web, and use our modification text generation language model (MTG-LLM) to describe the difference between the two captions. MTG-LLM is trained on a dataset of 715 triplet text annotations (Brooks, Holynski, and Efros 2023). The resulting triplet with the two corresponding videos (query $q$ and target video $v$) and the modification text ($t$) is therefore obtained fully automatically, allowing a scalable CoVR training data generation.
We apply our triplet generation approach to the WebVid2M dataset (Bain et al. 2021) which contains 2.5M Web-scraped video-caption pairs. This results in the WebVid-CoVR training dataset with 1.6M CoVR triplets. By virtue of its automatic generation procedure, WebVid-CoVR is inherently noisy. To efficiently train on such large-scale and noisy data, we use a contrastive loss (van den Oord, Li, and Vinyals 2018), adopting the HN-NCE variant from (Radenovic et al. 2023) to upsample the significance of hard negatives. We design a CoVR model based on the cross-modal BLIP (Li et al. 2022b) and use query scoring (Bain et al. 2022) to exploit information from multiple video frames. Training this model on WebVid-CoVR shows strong transferability to the CoIR task, in both zero-shot and finetuning settings, achieving state-of-the-art results on the standard CIRR (Liu et al. 2021) and FashionIQ (Wu et al. 2021) benchmarks. Finally, to foster research in CoVR, we repeat our generation procedure on a distinct subset of the WebVid10M dataset (Bain et al. 2021) and manually select correctly generated samples to constitute WebVid-CoVR-Test, a test set of 2,435 CoVR triplets. We find that our model achieves promising results on WebVid-CoVR-Test compared to standard baselines.
To summarize, our contributions are: (i) We propose a scalable approach to automatically generate composed visual retrieval training data. We apply this pipeline to the WebVid2M dataset and generate the WebVid-CoVR training dataset with 1.6M CoVR triplets. (ii) We show that training a CoVR model on WebVid-CoVR transfers well to the CoIR task, and achieves state-of-the-art results on the CIRR and FashionIQ benchmarks in the zero-shot and finetuning setups. (iii) We evaluate our model on WebVid-CoVR-Test, a new CoVR benchmark that we manually annotate. Our code, datasets, and models are available at https://imagine.enpc.fr/~ventural/covr.
2 Related Work
Dataset Type #Triplets #Visuals #Unique words Avg. text length Domain
[…]
Composed image retrieval (CoIR). CoIR (Vo et al. 2019) has been an active area of research in recent years (Guo et al. 2018; Vo et al. 2019; Wu et al. 2021; Baldrati et al. 2022; Liu et al. 2021; Delmas et al. 2022; Kim et al. 2021; Baldrati et al. 2023; Saito et al. 2023; Gu et al. 2023). Most methods designed for this problem use manually annotated image-text-image triplets for training (Wu et al. 2021; Liu et al. 2021; Baldrati et al. 2022; Delmas et al. 2022). Very recent works, such as Pic2Word (Saito et al. 2023) and SEARLE (Baldrati et al. 2023), explore zero-shot CoIR setups where no manually annotated CoIR triplet is used. These approaches build on CLIP (Radford et al. 2021) and train a mapping network using image-only data for text inversion so that they can be flexibly composed with text descriptions. Our approach is similar in that it avoids collecting manual triplets; however, we instead perform supervised training on automatically generated image-text-video triplets given only video-text pairs. We also differ from above works by focusing on the composed video retrieval (CoVR) task, as opposed to CoIR.
Datasets for composed image retrieval. CIRR (Liu et al. 2021) and Fashion-IQ (Wu et al. 2021) are the two most widely used CoIR benchmarks. Both are manually annotated, hence small scale (about 30k triplets, see Table 1) due to the high cost implied in collecting CoIR triplets. To scale up, two concurrent works proposed larger, automatically generated CoIR datasets: LaSCo (Levy et al. 2024) and SynthTriplets18M (Gu et al. 2023). However, these two datasets are currently not publicly available. The LaSCo dataset (Levy et al. 2024) is generated using the visual question answering annotations and the pairing between images and counterfactual images in the VQAv2 dataset (Antol et al. 2015). In detail, this dataset provides for each (image, question, answer) triplet a counterfactual triplet with the same question and different image and answer. In contrast, we do not rely on such expensive annotation schemes. SynthTriplets18M (Gu et al. 2023) uses the text-conditioned image editing framework InstructPix2Pix (Brooks, Holynski, and Efros 2023) to automatically generate CoIR data. Their edit text generation process is similar to ours, but our generation process differs in that we automatically mine similar videos from a dataset of video-text pairs to construct CoVR triplets instead of generating visual data. In experiments, we show the superiority of our triplet construction procedure as we achieve much higher CoIR results (e.g., 39% vs 19% zero-shot R@1 on CIRR while generating fewer data). Lastly, our WebVid-CoVR dataset is not limited to still images and considers videos, while standing out as the largest composed retrieval dataset in the natural domain, as depicted in Table 1.
Vision-language pretraining. Many strong multi-modal models have been pretrained on large datasets of image-caption pairs (Alayrac et al. 2022; Chen et al. 2020; Li et al. 2021, 2023, 2020b; Radford et al. 2021; Jia et al. 2021; Su et al. 2019; Schuhmann et al. 2022) or video-caption pairs (Akbari et al. 2021; Li et al. 2022a, 2020a; Miech et al. 2019, 2020; Xu et al. 2021; Sun et al. 2022; Xue et al. 2022a; Zhao et al. 2023). In contrast, we generate CoVR training data from video-caption pairs instead of directly training on them. Our data generation approach is also related to other generation approaches used for other tasks, e.g., action recognition (Nagrani et al. 2020), visual question answering (Yang et al. 2021), and visual dialog (Liu et al. 2023). However, unlike all these tasks, the CoVR task requires retrieving visual data.
Video retrieval. Text-to-video retrieval has received great attention over the last few years (Fang et al. 2021; Gao et al. 2021; Ge et al. 2022; Liu et al. 2022; Luo et al. 2021; Ma et al. 2022; Rasheed et al. 2023; Xue et al. 2022b; Yang, Bisk, and Gao 2021; Yao et al. 2022; Xu et al. 2021). We also make use of multiple video frames with query scoring similar to (Bain et al. 2022). However, different from these methods, we focus on composed video retrieval, where the query consists of both text and visual data.
3 Automatic Triplet Generation and Training
The goal of our composed video retrieval (CoVR) task is, given an input image $q$ and a modification text $t$, to retrieve a modified video $v$ in a large database of videos [^1] Note that $q$ could also be a video query, but in our main experiments we focus on an image query, and provide more results in the supplementary material (Section C.2) with video queries.. Our goal is to avoid the manual annotation of $(q,t,v$) triplets for training. Hence we automatically generate such triplets from Web-scraped video-caption pairs, as explained in Section 3.1 and illustrated in Figure 2. The resulting WebVid-CoVR dataset, together with its manually curated evaluation set, is presented in Section 3.2. Finally, we present how we train a CoVR model using WebVid-CoVR in Section 3.3.
3.1 Generating Composed Video Retrieval Triplets
Given a large (Web-scraped) dataset of video-caption pairs, we wish to automatically generate video-text-video CoVR triplets $(q,t,v$) where the text $t$ describes a modification to the visual query $q$. However, the dataset of video-caption pairs neither contains annotations of paired videos, nor modification text that describes their difference. Hence we propose a methodology to automatically mine paired videos and describe their difference, as described below. Note that for illustration, we take as an example the WebVid2M dataset (Bain et al. 2021) with 2.5M video-caption pairs, but this methodology could potentially be applied to other large datasets of video-text (or image-text) pairs.
Mining paired videos by pairing captions. In order to obtain video pairs that exhibit visual similarity while differing in certain aspects, we leverage their associated captions. The core idea is that videos with similar captions are likely to have similar visual content. Specifically, we consider captions that differ by a single word, excluding punctuation marks. For instance, the caption “Young woman smiling” is paired with “Old woman smiling” and “Young couple smiling”. In the 2M distinct captions from WebVid2M, this process allows us to identify a vast pool of 1.2M distinct caption pairs with 177k distinct captions, resulting in 3.1M paired videos. In the following, we describe further steps to filter the data into a smaller set.
Filtering caption pairs. We wish to automatically generate the modification text between paired videos using their (paired) captions. However, caption pairs with the same meaning are likely to result in meaningless differences. On the contrary, caption pairs that differ too much are likely to result in large visual differences that cannot be easily described. To address these issues, we filter out caption pairs that are too similar and too dissimilar. Specifically, we exclude caption pairs with CLIP text embedding similarity $\geq$ 0.96 (e.g., “Fit and happy young couple playing in the park” and “Fit and happy young couple play in the park”) and caption pairs with CLIP text embedding similarity $\leq$ 0.6 (e.g., Zebra on a white background” and Coins on a white background”). We also exclude pairs where the captions differ by a digit (which mostly consist of a date in practice), a word not part of the English dictionary, or by a rare word. Rare words are detected based on the zipfzipf frequency (Speer 2022). Finally, we remove templated captions such as “abstract of”, “concept of”, and “flag of” which are over-represented in WebVid2M. At the end of this filtering stage, we have 370k distinct caption pairs with 92k distinct captions, resulting in 1.2M paired videos that we will use to generate the modification text. Note that we can use these paired videos in both directions to generate triplets, as the source and target videos can be swapped.
Generating a modification text from paired captions. In order to generate a modification text between paired videos, we develop and apply a “modification text generation large language model” (MTG-LLM) to their corresponding paired captions. We describe the MTG-LLM inference process below and then explain its training details. The MTG-LLM takes as input two paired captions and generates a modification text that describes the difference between the two captions (see Figure 2). In detail, the generation is auto-regressive, i.e., we recursively sample from the token likelihood distribution conditioned on the previously generated tokens until an end-of-sentence token is reached. Examples of the input-output, and details about the prompt format, which involves concatenating the two captions with a delimiter, can be found in Section B.4 of the Appendix. We use top-k sampling (Fan, Lewis, and Dauphin 2018) for generating the tokens instead of maximum-likelihood methods such as beam search. Note that we only generate a single modification text per caption pair for computational efficiency, but the MTG-LLM could be used to generate multiple modification texts per caption pair which could serve as a data augmentation in future work.
We now describe the training details of the MTG-LLM. We start from a LLM pretrained with a next token prediction objective on a Web-scale text dataset, namely LLaMA (Touvron et al. 2023). We then finetune this LLM for the MTG task on a manually annotated text dataset. In particular, we repurpose the editing dataset from InstructPix2Pix (Brooks, Holynski, and Efros 2023), which provides a modification text and a target caption for 700 input captions. We augment this dataset with 15 annotations that cover additional cases. More details about the additional examples can be found in Section B.4 of the Appendix.
Filtering video pairs. We wish to avoid some modification texts being over-represented in the dataset as it could harm training. Hence, if there are more than 10 video pairs associated with the same pair of captions (therefore leading to the same modification text), we only select top 10 video pairs. As the CoVR task typically involves similar query-target video pairs, we choose pairs of videos with the highest visual similarity, as measured by the CLIP visual embedding similarity computed at the middle frame of the videos.
Figure 3: Examples of generated CoVR triplets in WebVid-CoVR: The middle frame of each video is shown with its corresponding caption, with the distinct word highlighted in bold. Additionally, the generated modification text is displayed on top of each pair of videos. The bottom example illustrates a noisy generated modification text, as ‘beautiful’ is subjective and both target and query videos can be considered as beautiful fields.
3.2 Our Resulting WebVid-CoVR Dataset
In the following, we describe the training and test partitions of our CoVR data. While our training set is automatically generated, our test set is manually verified.
WebVid-CoVR: a large-scale CoVR training dataset. By applying the previously described pipeline to the WebVid2M dataset (Bain et al. 2021), we generate WebVid-CoVR, a dataset containing 1.6M CoVR triplets, which is significantly larger than prior datasets (see Table 1). On average, a video lasts 16.8 seconds, a modification text contains 4.8 words, and one target video is associated with 12.7 triplets. We study the effect of the modification text length in Section C.6 of the Appendix. WebVid-CoVR is highly diverse with 131k distinct videos and 467k distinct modification texts. Examples of CoVR triplets from the WebVid-CoVR dataset are illustrated in Figure 3. These examples (along with additional ones included in Section D.3 of the Appendix) demonstrate the diversity present in WebVid-CoVR, highlighting a wide range of content and variations in the modification texts. However, it is important to acknowledge that some noise naturally exists in the dataset, as shown in the bottom example of Figure 3, where the text does not describe the difference between the two videos due to both videos describing beautiful fields. We provide further analysis such as removal of inappropriate content, and dataset statistics of WebVid-CoVR in Section A of the Appendix.
WebVid-CoVR-Test: a new CoVR evaluation benchmark. Due to the noise in WebVid-CoVR, we manually annotate a small test set, dubbed WebVid-CoVR-Test, for evaluation. For this, we first repeat the data generation procedure described in Section 3.1, but on a different corpus of video-caption pairs. Specifically, we consider video-caption pairs from the WebVid10M corpus (Bain et al. 2021) that are not included in the WebVid2M dataset, resulting in a pool of 8 million video-caption pairs. This ensures that other models using WebVid2M for pretraining have not been exposed to any of the test examples. In the video pairs filtering stage, for each pair of captions, we here only keep one pair of videos (the one with the highest visual similarity). This results in 163k candidate triplets that could be used for testing purposes. We randomly sample 7k triplets that we use for validation and randomly sample 3.2k other triplets that we manually annotate as described below.
We augment the 3.2k triplets by generating two additional modification texts with the MTG-LLM. The annotator reads the three generated modification texts, looks at three frames from the query and target videos, and either keeps the best modification text if at least one is valid or discards the sample. Through this meticulous annotation process, we ensure that the test set comprises high-quality and meaningful CoVR triplets. This results in a test set of 2.5k triplets, i.e., about 22% of the examples are considered as noisy and are discarded.
3.3 Training on WebVid-CoVR
Here, we describe our CoVR model architecture and how we train it on our WebVid-CoVR dataset.
CoVR-BLIP model architecture. Our model architecture builds on a pretrained image-text model, BLIP (Li et al. 2022b). BLIP is pretrained on a large dataset of image-caption pairs with three vision-language objectives: image-text contrastive learning, image-text matching, and image-conditioned language modeling. However, BLIP is not trained for composed visual retrieval with both visual and text inputs. Therefore we adapt BLIP to the CoIR/CoVR task as follows.
[…]
Training. In order to train on WebVid-CoVR, we use a contrastive learning approach (van den Oord, Li, and Vinyals 2018; Radenovic et al. 2023), as it has been shown to be effective to learn strong multi-modal representations from large-scale noisy data (Radford et al. 2021). We make the following design choices. First, we create a training batch by sampling distinct target videos; and for each target video, we randomly sample an associated image-text query pair. Iterating over videos ensures that the same target video appears only once in a batch and maximizes the number of different target videos that can be used as negatives in contrastive learning. We show the benefit of this approach in Section 4.4 (Table 6). Second, we employ HN-NCE (Radenovic et al. 2023) which increases the weight of most similar samples and uses as negatives all target videos $v_{j\in\mathcal{B}}$ in the batch $\mathcal{B}$. Formally, given a training batch $\mathcal{B}$ of triplets $(q_{i},t_{i},v_{i})$, we minimize the following loss:
$\mathcal{L}(\mathcal{B})=$ $-\sum_{i\in\mathcal{B}}\text{log}\left(\frac{e^{S_{i,i}/\tau}}{\alpha\cdot{e^{S_{i,i}/\tau}}+\sum_{j\neq i}{e^{S_{i,j}/\tau}w_{i,j}}}\right)$
$-\sum_{i\in\mathcal{B}}\text{log}\left(\frac{e^{S_{i,i}/\tau}}{\alpha\cdot{e^{S_{i,i}/\tau}}+\sum_{j\neq i}{e^{S_{j,i}/\tau}w_{j,i}}}\right)$
where $\alpha$ is set to 1, temperature $\tau$ is set to $0.07$, $S_{i,j}$ is the cosine similarity between the multi-modal embedding $f_{i}$ and the target video embedding $\hat{v}_{j}$, and $w_{i,j}$ is set as in (Radenovic et al. 2023) with $\beta=0.5$.
4 Experiments
We first describe the experimental protocol including the datasets, evaluation metrics, and implementation details (Section 4.1). We then present the results of CoVR on our new video benchmark (Section 4.2), as well as transfer results of CoIR on standard image benchmarks (Section 4.3). Finally, we provide ablations on our key components (Section 4.4).
4.1 Experimental Setup
Datasets. WebVid-CoVR is our proposed training CoVR dataset, and WebVid-CoVR-Test is our new CoVR benchmark, both presented in Section 3.2.
CIRR (Liu et al. 2021) is a manually annotated CoIR dataset that contains open-domain natural images from NLVR2 (Suhr et al. 2019), comprising 36.5k queries annotated on 19k images. CIRR includes two evaluation protocols: a standard one with the entire validation set as the search gallery, and a fine-grained subset, where the search space is a subgroup of six images similar to the query image (based on pretrained ResNet15 feature distance). The dataset is divided into training, validation, and testing splits with 28225/16742, 4181/2265 and 4148/2178 queries/images, respectively.
FashionIQ (Wu et al. 2021) is another CoIR dataset that contains images of fashion products, divided into three categories of Shirts, Dresses, and Tops/Tees. The query and target images were automatically paired based on title similarities (crawled from the web), and modification texts were then manually annotated. This dataset consists of 30k queries annotated on 40.5k different images. It is divided into training and validation splits with 18000/45429 and 6016/15415 queries/images, respectively.
Evaluation metrics. Following standard evaluation protocols (Liu et al. 2021), we report the video retrieval recall at rank 1, 5, 10, and 50. Recall at rank k (R@k) quantifies the number of times the correct video is among the top k results. MeanR denotes the average of R@1, R@5, R@10, and R@50. Higher recall means better performance.
Input WebVid-CoVR-Test
modalities Fusion Backbone #frames R@1 R@5 R@10 R@50
Random - - - - 00.08 00.23 00.35 01.76
Text - BLIP - 19.68 37.09 45.85 65.14
Not finetuned on Visual - BLIP 15 34.90 59.23 68.04 85.95
WebVid-CoVR Visual + Text Avg CLIP 15 44.37 69.13 77.62 93.00
Visual + Text Avg BLIP 15 45.46 70.46 79.54 93.27
Text - BLIP - 23.67 45.89 55.13 77.03
Visual - BLIP 15 38.89 64.98 74.02 92.06
Finetuned on Visual + Text MLP CLIP 1 50.55 77.11 85.05 96.60
WebVid-CoVR Visual + Text MLP BLIP 1 50.63 74.80 83.37 95.54
Visual + Text CA BLIP 1 51.80 78.29 85.84 97.07
Visual + Text CA BLIP 15 53.13 79.93 86.85 97.69
Table 2: Benchmarking on the WebVid-CoVR-Test set: We observe that using both the visual and text input modalities performs better than individual modalities alone, both with/without finetuning on WebVid-CoVR (shown at the top/bottom of the table, respectively). When using pretraining models without finetuning, we apply average fusion (Avg) for the embeddings. BLIP performs slightly better than CLIP on this benchmark. Finetuning on WebVid-CoVR brings significant benefits. In this case, fusing with the pretrained cross-attention (CA) from BLIP is more effective than training a randomly-initialized MLP fusion as done in (Baldrati et al. 2022). Moreover, using multiple frames to embed the target video brings further improvements over using the middle frame. The first row represents the random performance.
Implementation details. For our MTG-LLM, we use LLaMA 7B model (Touvron et al. 2023) that we finetune for one epoch with an initial learning rate of $3\mathrm{e}{-5}$. For our CoVR model, we use the BLIP model with ViT-L (Dosovitskiy et al. 2021) at 384 pixels finetuned for text-image retrieval on COCO and freeze the ViT for computational efficiency. We train our CoVR model on WebVid-CoVR for 4 epochs with a batch size of 2048 and an initial learning rate of $1\mathrm{e}{-5}$. To finetune on CIRR/FashionIQ, we train for 6 epochs with a batch size of 2048/1024 and an initial learning rate of $1\mathrm{e}{-4}$. We set hyperparameters based on the validation curve of WebVid-CoVR. Experiments are conducted on 4 NVIDIA A100-SXM4-80GB GPUs. More details are included in Section B of the Appendix.
4.2 Composed Video Retrieval Results
We provide a number of baselines for our new benchmark on WebVid-CoVR-Test. Table 2 summarizes these CoVR results. We first report the random chance performance in the first row. The rest of the table is split into two. The top block uses existing pretrained text and image encoders from BLIP (Li et al. 2022b) or CLIP (Radford et al. 2021) backbones without any finetuning. Models in the bottom block are finetuned on WebVid-CoVR. We report results with the composed query, as well as with the individual modalities. For combining modalities, we experiment with the simple average fusion baseline (Avg) when using frozen embeddings, and fusion with a randomly-initialized MLP or BLIP-pretrained cross-attention (CA) layers when finetuning. Note that the MLP fusion baseline is similar to Combiner (Baldrati et al. 2022) that concatenates the image and text embeddings from CLIP (or BLIP in CASE (Levy et al. 2024)), and is referred to as late fusion by CASE. For finetuning individual modalities, we train and test either with text-only query using the modification text, or with the visual-only image query. Finally, we experiment with using only the middle frame embedding or the weighted average of target video frame embeddings as explained in Section 3.3 (with the exception that visual-only experiments use equal weights due to not having access to the modification text for computing the scores).
Mode Pretraining CIRR FashionIQ
Method Data R@1 R@5 R@10 R@50 R@10 R@50
Train TIRG+LastConv (Vo ’19) $\dagger$ - 11.04 35.68 51.27 83.29 - -
TIRG (Vo ’19$\dagger$ - 14.61 48.37 64.08 90.03 - -
JVSM (Chen ‘20) - - - - - 11.90 26.60
TRACE w/BER (Jandial ‘20) - - - - - 22.57 46.19
VAL w/GloVe (Chen ‘20) - - - - - 24.15 46.61
CurlingNet (Yu ‘20$\ddagger$ - - - - - 25.90 51.01
RTIC-GCN (Shin ‘21$\ddagger$ - - - - - 28.18 53.09
CoSMo (Lee ‘21) - - - - - 26.58 52.31
DCNet (Kim ‘21) - - - - - 27.78 53.89
MAAF (Doods ‘20$\dagger$ - 10.31 33.03 48.30 80.06 24.30 48.80
MAAF-BERT (Doods ‘20$\dagger$ - 10.12 33.10 48.01 80.57 - -
MAAF-IT (Doods ‘20$\dagger$ - 09.90 32.86 48.83 80.27 - -
MAAF-RP (Doods ‘20$\dagger$ - 10.22 33.32 48.68 81.84 - -
CIRPLANT (Liu ‘21$\dagger$ - 19.55 52.55 68.39 92.38 18.87 41.53
SAC w/BERT (Jandial ‘22) - - - - - 29.08 54.70
FashionVLP (Goenka ‘22) - - - - - 34.27 62.51
ARTEMIS (Delmas ‘22) - 16.96 46.10 61.31 87.73 26.05 50.29
LF-CLIP (Baldrati ‘22) - - - - - 35.39 59.03
LF-BLIP (Levy 2023) - 20.89 48.07 61.16 83.71 25.75 43.98
Combiner (Baldrati ‘22) - 33.59 65.35 77.35 95.21 - -
CompoDiff (Gu ‘23) ST18M 22.35 54.36 73.41 91.77 40.98 62.23
CASE (Levy ‘23) - 48.00 79.11 87.25 97.57 - -
CASE (Levy ‘23) LaSCo 48.68 79.98 88.51 97.49 48.79 70.68
CASE (Levy ‘23) LaSCo+COCO 49.35 80.02 88.75 97.47 - -
CoVR-BLIP - 49.16 79.76 88.65 97.49 49.02 70.24
CoVR-BLIP (Ours) WebVid-CoVR 50.41 80.96 89.35 97.64 49.40 70.98
[…]
CompoDiff (Gu ‘23) ST18M 19.37 53.81 72.02 90.85
Pic2Word (Saito ‘23) CC3M 23.90 51.70 65.30 87.80 24.70 43.70
CASE (Levy ‘23) LaSCo 30.89 60.75 73.88 92.84 - -
CASE (Levy ‘23) LaSCo+COCO 35.40 65.78 78.53 94.63 - -
CoVR-BLIP - 19.95 41.98 52.31 73.54 16.00 32.77
CoVR-BLIP (Ours) WebVid-CoVR 39.28 68.22 78.94 94.65 27.70 44.63
Table 3: State-of-the-art comparison on the CIRR test set and FashionIQ validation set: Our model benefits from training on WebVid-CoVR in both the zero-shot and finetuning setting. For full metrics, see Section C.8 of the Appendix. $\dagger$ denotes results reported by (Liu et al. 2021). For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. CC3M denotes Conceptual Captions 3M (Sharma et al. 2018), ST18M denotes  (Gu et al. 2023), and COCO denotes (Lin et al. 2014).
We make several conclusions. (i) Combining both visual and text modalities yields better performance than the models with individual modalities. This result highlights that our new CoVR benchmark requires paying attention to both modalities. (ii) Visual-only outperforms text-only suggesting that the video pairs automatically mined through their caption similarity indeed exhibits visual similarity, and that the image captures the target video better than the modification text. (iii) Finetuning on WebVid-CoVR obtains substantial improvements over using pretrained and frozen embeddings. (iv) When finetuning, fusion with BLIP cross-attention (CA) performs better than the MLP fusion. (v) Results with the BLIP backbone are marginally higher than those with CLIP. (vi) Using $N=15$ target video frames further boosts the performance over using only the middle frame. We further analyse the effect of the number of frames in Section C.7 of the Appendix.
4.3 Transfer Learning to CoIR
While our focus is video retrieval, we also experiment with transferring our CoVR models to image retrieval tasks on standard CoIR benchmarks. We define zero-shot CoIR as not using any manually annotated CoIR triplet for training. We perform zero-shot CoIR by directly applying our model trained on our automatically generated WebVid-CoVR dataset to CoIR tasks and also explore finetuning our model on the training set of the downstream benchmark.
Table 3 report results on CIRR and Fashion-IQ datasets. These results show that our model highly benefits from training on WebVid-CoVR, especially in the zero-shot setting, on both datasets. In addition, our model achieves state-of-the-art performance in both zero-shot and finetuning settings on the CIRR and FashionIQ datasets.
4.4 Ablation Studies
We now ablate the importance of several key aspects of our method by training only on WebVid-CoVR.
Importance of data scale. In Table 4, we evaluate the effect of the number of video-caption pairs used as a seed for our triplet generation pipeline. We construct subsets of videos such that larger ones include smaller ones, and only keep triplets that contain the sampled videos for training. We find that results steadily increase when using more videos, demonstrating that our method largely benefits from scaling the size of the seed dataset of video-captions. We also observe the importance of the filtering techniques described in Section 3.1, as the model trained on unfiltered data underperforms.
Initial Generated W-C-T CIRR FashionIQ
#videos #triplets Fil. MeanR MeanR MeanR
0 - - 36.80 46.95 24.39
200k 11k 62.20 67.01 37.04
500k 66k 76.12 69.92 36.98
1M 269k 77.89 70.03 36.47
2.5M 1.6M 79.40 70.27 36.17
2.5M 3.6M 78.92 69.25 32.31
Table 4: Data size: We measure the importance of the number of videos used for data generation and of filtering (Fil. column) the generated data, by evaluating on WebVid-CoVR-Test (W-C-T), CIRR, and FashionIQ. All models are trained for the same number of iterations on the generated data.
Modification text generation. We use a large language model finetuned for modification text generation (MTG-LLM) as explained in Section 3.1. We here compare this solution to a simple rule-based baseline that uses several templates to generate the modification text given the two captions that differ by one word. Specifically, the modification text is based on the two different words from the captions. We generate templates that use these words and choose one at random during training. These templates include variations such as ‘Remove $\mathtt{txt\_diff_{1}}$” and "Change $\mathtt{txt\_diff_{1}}$ for $\mathtt{txt\_diff_{2}}$. A full list of all the templates can be seen in Section B.3 of the Appendix. Additionally, we investigate the possibility of paraphrasing the rule-based modification texts using GPT-3.5-turbo from OpenAI (Brown et al. 2020) as a source of augmentation, by prompting “Paraphrase the following sentence: {Rule-base modification text}. In preliminary analysis, we qualitatively observed that LLaMA (Touvron et al. 2023) and LLaMA 2 (Touvron et al. 2023) alternatives were overly verbose when used for paraphrasing; however, GPT-3.5 outputs were satisfactory.
In Table 5, we show that our MTG-LLM generates better modification texts than the rule-based baseline, by evaluating the results of the model trained on the generated data. Paraphrasing the rule-based examples significantly boosts the performance (from 41 to 52 R@1), while still being worse than our MTG-LLM, especially on the CIRR benchmark. Note that the paraphrasing comes with the cost of running an expensive LLM (\$43 cost for this experiment for 1 paraphrasing per modification text on the entire dataset). On the other hand, our MTG-LLM finetuning only requires 715 text examples. Qualitative examples comparing MTG-LLM and rule-based are provided in Table A.11 of the Appendix.
WebVid-CoVR-T. CIRR
Model R@1 MeanR R@1 MeanR
Rule-based 41.39 70.36 16.31 47.84
Rule-based paraph. 52.39 78.75 34.53 66.54
MTG-LLM 53.13 79.40 39.28 70.27
Table 5: Modification text generation: We compare our MTG-LLM (LLaMA 7B parameters) against both a rule-based MTG baseline and a paraphrased rule-based MTG baseline (using GPT-3.5-turbo from OpenAI). We observe important gains in the downstream performance of the model trained on the generated data.
WebVid-CoVR-T. CIRR
Iteration HN-NCE R@1 MeanR R@1 MeanR
Triplets 49.80 78.15 35.66 67.15
Videos 49.02 76.62 35.93 67.72
Videos 53.13 79.40 39.28 70.27
Table 6: Training strategies: Iterating on batches of distinct target videos (instead of triplets) and up-sampling hard negatives (HN-NCE (Radenovic et al. 2023)) both benefit the CoVR/CoIR performance.
Training strategies. In Table 6, we first show the benefit on WebVid-CoVR of training by iterating on target videos instead of CoVR triplets. This is to avoid having the same target video appearing multiple times in a training batch, hence increasing the number of correct negatives that are used in the contrastive loss. Furthermore, up-sampling hard negatives adopting the HN-NCE loss formulation from (Radenovic et al. 2023) also slightly benefits the performance.
5 Conclusions and Limitations
In this work, we studied the new task of CoVR by proposing a simple yet effective methodology to create automatic training data. Our results on several benchmarks (including our manually curated video benchmark, as well as existing image benchmarks) suggest that, while noisy, such an automated and scalable approach can provide effective CoVR model training. One potential limitation of our method is that the generated modification text may not depict some visible changes due to not considering the image pair, but only their captions. Moreover, our modification text is suboptimal due to only inputting one-word difference caption pairs (i.e., focusing only on one change, and not considering multi-word differences). For example, the following modification with multiple changes from the CIRR dataset would not exist in our data “close up of a similar dog, but it is swimming on its own with a tennis ball in its mouth”. Future work can incorporate visually-grounded modification generation and multiple modifications between query and target video pairs.
Ethics statement
Our model constitutes a generic multi-modal search tool, but is not intended for a specific application. While there are helpful use cases such as online shopping, traveling, and personal development (i.e., how-to), there may be potential privacy and harmful risks when training the model on datasets with inappropriate content. The risks include surveillance applications such as searching for a specific person, and looking up violent and graphic videos. For our WebVid-CoVR data release, we refer to Section A for further analysis about removal of inappropriate content. We note that our dataset users must also adhere to the terms of use stipulated by WebVid (Bain et al. 2021).
Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014223 made by GENCI. The authors would like to acknowledge the research gift from Google, the ANR project CorVis ANR-21-CE23-0003-01, Antoine Yang’s Google PhD fellowship, and thank Mathis Petrovich, Nicolas Dufour, Charles Raude, and Andrea Blazquez for their helpful feedback.
APPENDIX
This document provides dataset statistics (Section A), implementation details (Section B), additional experiments (Section C), and qualitative examples (Section D). We also provide the code and dataset together with a datasheet, and an illustrative video on our project page at imagine.enpc.fr/~ventural/covr.
Appendix A Dataset statistics and analysis
In this section, we provide analysis on our WebVid-CoVR. A detailed datasheet can be found as a separate file.
Filtering inappropriate content and vulgar language. We take several measures to detect semi-automatically any inappropriate content, and remove such instances from our dataset. To achieve this, we use a combination of tools (such as negative sentiment and profanity detectors) and apply them on modification texts and video captions.
We conduct a sentiment analysis on the modification texts using the TextBlob library (Loria 2013) to identify instances of negative sentiment. We find that less than 0.5% of the dataset (about 2k instances) exhibits negative sentiment. Upon manual review, we identify false positives in this categorization, including examples such as “make it an evil pumpkin” or “Change him into a frustrated businessman”. The instances detected as negative sentiment are reviewed and 260 of them are removed from the dataset. We ensure that the dataset does not include any videos marked for mature content, by checking the metadata of WebVid (Bain et al. 2021) provided by (Lab 2023). Finally, using the better-profanity library (Thanh 2018), we identify approximately 2k video captions that are marked for profanity. Upon manual inspection, we find that there were a large number of videos displaying computer-generated visuals with those words. We also notice false positives (e.g., misinterpretation due to context), such as the animal cock being incorrectly identified as profanity. The videos detected to contain profanity in their captions are reviewed and excluded from the dataset.
Distribution of caption and video embedding similarities. As explained in Section 3.1 of the main paper, we filter caption pairs with CLIP text embedding similarity $\geq$ 0.96 and caption pairs with CLIP text embedding similarity $\leq$ 0.6, and for each caption pair, we choose the 10 video pairs with the highest CLIP visual similarity computed at the middle frame of the videos. We also note that our cosine similarities are normalized between [0, 1]. Here, we further show the distribution of text embedding similarity in caption pairs and visual embedding similarity in video pairs in Figure A.1. The distribution of video similarity scores exhibits two distinct peaks. The first peak corresponds to a score of approximately 0.7 and includes video pairs that are significantly dissimilar. The second peak corresponds to a score close to 1.0 and represents video pairs with highly similar visual content.
Number of words in modification texts. Figure A.2 further provides the histogram of the number of words in the generated modification text. We observe that the majority of texts contain 3-8 words.
Number of triplets per target video. In Section 3.2 of the main paper, we provided several statistics about our WebVid-CoVR dataset, e.g., on average, a target video is associated with 12.7 triplets. However, in Figure A.3, when visualizing the distribution of triplets associated with each target video, we see that the histogram reveals that the majority of target videos are associated to only 1 or 2 triplets. The histogram exhibits a long tail, i.e., a small subset of target videos have a considerably larger number of triplets associated. These videos have captions such as “Mountain landscape”, “Water stream”, and “Water river”, leading to numerous one-word difference captions associated with them.
Video categories. We plot the distribution of video categories in Figure A.4. These categories are found using the WebVid metadata provided by (Lab 2023). We find 50% of WebVid-CoVR videos in this metadata collection. Note more than one category can be associated with a single video (e.g., Nature and Animals/Wildlife for a video of a fish in the ocean).
Distribution of part-of-speech (POS) tags. We conducted POS tagging on the modification texts within the WebVid-CoVR dataset to analyze their distribution. The resulting analysis reveals the average counts of different parts of speech per modification text, including Nouns, Verbs, Pronouns, Adjectives, and Adverbs. We plot the distribution in Figure A.4, and see that, on average, a modification text contains 1.6 nouns and 1.1 verbs, emphasizing the prevalent use of nouns and verbs in the dataset’s modifications. The most frequently encountered words within each category’s top 3 are as follows: Noun: symbol, water, forest. Verb: make, turn, change. Pronoun: it, them, her. Adjective: green, more, black. Adverb: instead, more, then. We also include a visualization of the verb-noun frequency heatmap in Figure A.6, which provides insights into the distribution of verb-noun count combinations across modification texts in our dataset. From the heatmap, we observe that over 60% of the sentences exhibit a pattern of having one verb paired with one or two nouns.
We also conducted an analysis using POS tagging on the video captions. Figure A.7 visually illustrates the transition of POS tags across the difference words in Caption 1 and Caption 2. We observe a predominant pattern of noun-to-noun changes in our caption pairs.
Source of noise. As mentioned in Section 3.2 of the main paper, about 22% of the automatic collection can be considered as noisy, because this was the percentage of discarded triplets when manually curating the WebVid-CoVR test set. We expect a similar noise ratio in the training set. To inspect the noise in detail, we manually went over the triplet examples that were marked as unsuitable (therefore discarded) when annotating the test set. We marked whether the reason for discarding falls within any of the following categories, and computed the following percentages (normalized by the number of discarded triplets).
  • 35%: The generated modification text does not describe the visual difference. Primarily attributed to either the quality of the video captions or the output generated by the MTG-LLM.
  • 28%: Paired videos are visually too similar.
  • 15%: Paired videos are visually too different.
  • 13%: At least one of the videos is difficult to understand/low quality.
  • 9%: Captions are too similar (e.g., one-word difference does not change the meaning: “On the chairlift” and “Ride the chairlift”).
While the first category of errors is the largest, it is important to also note that our strict standards for the test set necessitated the discarding of many triplets that could potentially be useful for training.
Figure A.1: Text/video similarity of the caption/video pairs: Distribution of text similarity scores between caption pairs $(caption_{1},caption_{2})$ (left) and video similarity scores between video pairs $(video_{1},video_{2})$ (right), using CLIP embeddings and cosine similarity.
Figure A.2: Histogram of the number of words in the generated modification text: Most modification texts have between 3 and 8 words.
Figure A.3: Distribution of number of triplets per target video: We display the histogram depicting the number of triplets associated with each target video in the WebVid-CoVR dataset. Most target videos have 1 or 2 triplets and certain videos exhibit a high number of triplets (zoomed in to the tail on the right plot), e.g., some target videos are present in over 300 triplets, highlighting the variability in modification texts.
Figure A.4: Distribution of video categories: We plot the distribution of categories for videos in WebVid-CoVR, as provided by (Lab 2023) as WebVid metadata. Note that 50% of our WebVid-CoVR videos are present in this metadata collection. Looking at the distribution, we observe that around 40% and 20% of WebVid-CoVR are videos of Nature and People, respectively.
Figure A.5: Distribution of parts of speech in modification texts: Distribution of nouns, verbs, pronouns, adjectives, and adverbs in the modification text using part-of-speech (POS) tagging. On average, there are more than one noun and one verb per modification text.
Figure A.6: Verb-noun heatmap: This heatmap illustrates the percentage of modification texts containing specific combinations of verbs and nouns. Each cell represents the frequency of a particular verb-noun combination, and the values are presented as percentages. The color intensity indicates the relative frequency of occurrence. We observe that over 60% of the sentences exhibit a pattern of having one verb paired with one or two nouns.
Figure A.7: Transition of POS tags across the difference words between the two captions: The visualization primarily focuses on nouns, adjectives, and verbs, which constitute a significant proportion of modifications at 87% (comprising 65% nouns, 13% adjectives, and 9% verbs). The remaining words fall into categories where the POS tagger was unable to classify the word (12%) or adverbs (<1%).
Appendix B Implementation details
We describe the dataset generation computation time (Section B.1), further training details (Section B.2), provide the templates we use for our rule-based baseline (Section B.3), and details about our MTG-LLM finetuning and inference (Section B.4).
B.1 Dataset generation computation time
We outline the detailed computation time for each step of the dataset generation. The computation times below are obtained using a single NVIDIA RTX A6000, but it is important to note that most of the processes can be parallelized, which would significantly reduce the wallclock time required. In practice, we used 2 GPUs.
  • Text embedding extraction: We extracted text embeddings from 2 million distinct captions out of a total of 2.4 million video-caption pairs. This process completed in less than 2 hours.
  • Caption similarity search: To identify captions with one-word differences, we employed the faiss library (Johnson, Douze, and Jégou 2019) to select the 100 closest captions, avoiding the need to compare each caption against the entire set of 2 million captions. This optimization significantly reduced the search time, resulting in 2.5 hours.
  • Text similarity filtering: Thanks to the precomputed text embeddings, the text similarity filtering step incurred no additional time overhead. All the text filtering processes were completed in less than 5 minutes, even on a large pool of 1.2 million captions.
  • Video similarity computation: To filter by video similarity, we extracted the middle frame from approximately 135,000 videos and computed CLIP embeddings. This step takes approximately 3 hours.
  • MTG-LLM model finetuning: Finetuning for 715 examples takes less than 10 minutes. Note that the time required to finetune the MTG-LLM model is independent of the number of CoVR triplets we generate.
  • Modification text generation: This is the most time-consuming stage of the pipeline. It takes around 24 hours to process the 1.6 million caption pairs.
B.2 Training details
Here, we provide implementation details in addition to Section 4.1 of the main paper. In terms of the optimization algorithm, we utilize AdamW (Loshchilov and Hutter 2019). For our MTG-LLM, we finetune for one epoch with a batch size of 128 and a learning of $3\mathrm{e}{-5}$ that is warmed up linearly for the first $100$ steps and then kept constant. For our CoVR model, keeping the visual backbone frozen largely improves the efficiency of the training process: an epoch on the CIRR dataset takes 4 minutes with a frozen backbone and 25 minutes with a finetuned backbone, while leading to similar performance. During the training process, we employ several image data augmentations. These transformations include a random resized crop, where the input image is resized to a resolution of $384\times 384$. Additionally, we apply a random horizontal flip and random adjustments to contrast, brightness, sharpness, translation, and rotation. We use a weight decay of 0.05 and an initial learning rate of $1\mathrm{e}{-5}$ that is decayed to 0 following a cosine schedule over 10 epochs.
B.3 List of rule-based templates
In the ablation studies (Section 4.4 of the main paper), we introduced a rule-based MTG baseline. Here, in Table A.1, we show the templates used for the rules. We refer to Section D.2 (Table A.11) for qualitative comparison with our finetuned MTG-LLM.
Table A.1: Rule-based templates: For our rule-based MTG baseline, we randomly choose one of the below templates during training.
Remove $\mathtt{txt\_diff_{1}}$
Take out $\mathtt{txt\_diff_{1}}$ and add $\mathtt{txt\_diff_{2}}$
Change $\mathtt{txt\_diff_{1}}$ for $\mathtt{txt\_diff_{2}}$
Replace $\mathtt{txt\_diff_{1}}$ with $\mathtt{txt\_diff_{2}}$
Replace $\mathtt{txt\_diff_{1}}$ by $\mathtt{txt\_diff_{2}}$
Replace $\mathtt{txt\_diff_{1}}$ with $\mathtt{txt\_diff_{2}}$
Make the $\mathtt{txt\_diff_{1}}$ into $\mathtt{txt\_diff_{2}}$
Add $\mathtt{txt\_diff_{2}}$
Change it to $\mathtt{txt\_diff_{2}}$
B.4 Generating a modification text from paired captions with MTG-LLM
[…]
For instance, the model takes as input:
[…]
and is trained to generate the response:
[…]
At inference, we simply leave the response empty, and let the model autoregressively generate a modification text.
As mentioned in Section 3.1 of the main paper, we add 15 manually prepared text triplets to the existing 700 text triplets from (Brooks, Holynski, and Efros 2023) used for training. The motivation is to address specific CoVR cases not present in the original set of triplets, such as “remove clouds and reveal only sky” given input captions “Clouds timelapse” and “Sky timelapse”. We show these 15 samples in Table A.2.
Table A.2: Added examples to the MTG-LLM training: We add the below 15 examples to the set of 700 text triplets from (Brooks, Holynski, and Efros 2023).
Caption1 Clouds in the sky
Caption2 Airplane in the sky
Target output Add an airplane
Caption1 Woman with the tablet computer sitting in the city.
Caption2 Woman with tablet computer sitting in the park.
Target output In the park
Caption1 Walking swan
Caption2 White swan
Target output Change color to white
Caption1 Child playing on beach, sea waves view, girl spinning on coastline in summer 4k
Caption2 Child playing on beach, sea waves view, girl running on coastline in summer 4k
Target output Make her spin
Caption1 Aerial view of forest
Caption2 Aerial view autumn forest
Target output Change season to autumn
Caption1 Palm tree in the wind
Caption2 Palm trees in the wind
Target output Add more palm trees
Caption1 Schoolgirl talking on the phone
Caption2 Girl talking on the phone
Target output Make her older
Caption1 Clouds timelapse
Caption2 Sky timelapse
Target output remove clouds and reveal only sky
Caption1 Aerial view of a sailboat anchored in the mediterranean sea, vathi, greece.
Caption2 Aerial view of two sailboat anchored in the mediterranean sea, vathi, greece.
Target output Add one sailboat
Caption1 France flag waving in the wind. realistic flag background. looped animation background.
Caption2 Italian flag waving in the wind. realistic flag background. looped animation background.
Target output Swap the flag for an italian one
Caption1 Woman jogging with her dog in the park
Caption2 Woman playing with her dog in the park.
Target output Stop jogging and make them play
Caption1 Oil Painting Reproductions of by humans william-glackens
Caption2 Oil Painting Reproductions of zombies by william-glackens
Target output Replace the humans with zombies
Caption1 The girl who loved the sea by banafria
Caption2 The girl, wearing a hat, who loved the sea by banafria
Target output Put a hat on her
Caption1 famous painting Paris, a Rainy Day of Gustave Caillebotte
Caption2 famous painting Paris, a Sunny Day of Gustave Caillebotte
Target output Change it to more pleasant weather
Caption1 Bee on purple flower
Caption2 Bee on a flower
Target output Change color of the flower
The caption pairs from Table A.2 originate from failure cases of the initial iteration of our MTG-LLM on WebVid-CoVR. We manually corrected some of the failures by typing modification texts without looking at the corresponding videos, i.e., in a way we would want a text-only model to behave, by focusing on the changed words. We note that some cases are ambiguous, especially when the captions are not long or precise enough. For example, ‘walking swan’ and ‘white swan’ may not necessarily result in ‘change color to white’, but there is no way to know without looking at the visual pair, which itself is an interesting area for future work.
Appendix C Additional experiments
We provide additional experiments, reporting CoVR results obtained by training on data generated with prompting (i.e., without finetuning) the LLM (Section C.1), results when changing the visual query from an image to a video (Section C.2), varying the pretrained BLIP model (Section C.3), incorporating visual similarity between videos (Section C.4), results when filtering dynamic or static videos (Section C.5) effect of the modification text length (Section C.6), optimal number of frames (Section C.7), and full CoIR metrics (Section C.8).
C.1 Prompting versus finetuning the MTG-LLM
Here, we justify why we finetuned Llama as opposed to simply prompting it without any training. For prompting, we prepend few-shot examples of pairs of captions and desired generated texts, before adding the two captions in question. In particular, we use the following sentence:
[…]
Then, we concatenate our two captions for which we wish to generate a modification text. Table A.3 shows that finetuning the MTG-LLM for generating the training data is much more effective than prompting it without finetuning, as measured by CoVR performance on WebVid-CoVR-Test and CoIR performance on CIRR. This is also consistent with our qualitative observations: we found that the LLM struggles to perform the modification text generation without finetuning (see Table A.11 in the next section).
Table A.3: Prompting versus finetuning LLM: We compare our finetuned model (MTG-LLM) to a prompting baseline (see Section C.1) and observe important gains in the downstream performance of the model trained on the generated data.
WebVid-CoVR-Test CIRR
Model R@1 R@5 R@10 R@50 R@1 R@5 R@10 R@50
Prompting 51.33 76.68 85.13 96.71 35.59 64.67 76.19 92.87
Finetuning 53.13 79.93 86.85 97.69 39.28 68.22 78.94 94.65
C.2 Video query for CoVR
As noted in Section 3 of the main paper, we focus on image queries in this paper. This was because querying with an image has arguably more applications for realistic search scenarios. Here, we explore the setup of using a video as the visual query instead of an image query. We can do this since our dataset consists of video-text-video triplets. To encode a query video, we sample 5 equally-spaced frames and compute visual embeddings for each frame using the BLIP image encoder. We then average the per-frame embeddings and forward it through the BLIP cross-attention layers to obtain a multimodal query embedding $f(q,t)$. Note that we keep the target video representation fixed to 15 frames with weighted embedding averaging as described in Section 3.3 of the main paper. As seen in Table A.4, using 5 query frames leads to similar performance to using the middle frame.
Table A.4: Querying with a video: We report results on WebVid-CoVR-Test by using multiple frames from the query video. Recall that the rest of the paper investigates the setup where the middle video frame is used as an image query. To keep the computational complexity low, we only use 5 query video frames (uniformly sampled throughout the video). The number of target video frames remains unchanged as 15. The performance is similar to the image query setup, with marginal increase.
Visual query R@1 R@5 R@10 R@50
Image (middle frame) 53.17 79.93 86.85 97.69
Video (5 uniform frames) 53.91 79.85 87.09 97.42
C.3 Variants of pretrained BLIP backbones
All experiments in this paper are performed with the BLIP model (Li et al. 2022b) finetuned on COCO (Lin et al. 2014). Here, we include experiments when changing this backbone with other pretrained BLIP variants. Specifically, we use the BLIP model without COCO finetuning (BLIP base), and the BLIP model finetuned on Flickr30k (Plummer et al. 2015). For this experiment (as in the last row of Table 2 of the main paper), we use pretrained cross-attention layers of BLIP as our multimodal combined representation, and finetune them on WebVid-CoVR with 15-frame target video embeddings. In Table A.5, we observe that the BLIP model finetuned on COCO has the highest performance.
Table A.5: Variants of pretrained BLIP backbones: We compare the BLIP model without finetuning (base), BLIP finetuned on Flickr30k, and BLIP finetuned on COCO (the one used in the rest of the paper) (Li et al. 2022b). For this experiment, we finetune the models on WebVid-CoVR using the cross-attention layers of BLIP as the fusion method, and 15 frames for the target video as in the last row of Table 2.
Backbone R@1 R@5 R@10 R@50
BLIP Base 50.74 78.91 86.23 97.34
BLIP Flickr30k 52.50 79.46 86.70 97.77
BLIP COCO 53.13 79.93 86.85 97.69
C.4 Incorporating visual similarity between videos
When constructing the WebVid-CoVR dataset, we rely solely on caption similarity to identify similar video pairs for generating triplets, as discussed in Section 3.1.Here we provide an additional analysis on the effect of incorporating visual similarity in a similar fashion as we did with the caption similarity.
Specifically, we train variants of our model where we filter the WebVid-CoVR training triplets to only keep those whose video pair similarity is above a certain threshold. We measure the similarity using CLIP features on the middle frames, as in Section 3.1.The results in Table A.6 demonstrate that increasing the visual similarity threshold consistently decreases the downstream performance, while also discarding a large portion of the training data.
Table A.6: Effect of visual similarity: We observe worse performance on CoIR zero-shot benchmarks as we increase the visual similarity threshold in our training data. We train each model for the same number of iterations.
Threshold Data (%) CIRR R@1 FashionIQ R@10 mean
0.00 (None) 100% 39.28 27.70
0.55 092% 38.67 26.81
0.65 071% 37.54 25.77
0.70 055% 35.95 23.72
This suggests that relying solely on caption similarity is reasonable, and that incorporating visual similarity more strictly can be detrimental. A potential reason is that enforcing visual similarity constraints could bias the model to ignore the input text modification. Overall, our results indicate that the automatically mined video pairs with caption similarity already exhibit sufficient visual consistency for training the CoVR task.
C.5 Dynamic vs static content
The videos in our WebVid-CoVR dataset contain both dynamic sequences with motion, as well as static content. To analyze the prevalence, we compute the optical flow using the Gunnar Farneback’s algorithm (Farnebäck 2003) and empirically choose a magnitude threshold of 1 to distinguish between videos with static and dynamic elements. The magnitude value is obtained by averaging the Euclidean norms of motion vectors in both horizontal and vertical directions across the computed video frames. We identify that around 25% of the triplets contain static target videos, which represents approximately 21% of the overall target videos.
As a further experiment, we train models omitting either only the static portion or only the dynamic portions. The results in Table A.7 show that training on both dynamic and static triplets is beneficial, with a minor decrease in performance if we omit these static videos during training (while maintaining the same iteration count). This may be because image training data can still be complementary to video training (Bain et al. 2021).
Overall, we demonstrate that WebVid-CoVR contains both static and dynamic videos, hence posing an advantage over image datasets by providing more diversity.
Table A.7: Performance on WebVid-CoVR-Test when training on dynamic vs static video triplets: We see that training on both static and dynamic videos is beneficial.
Data (%) R@1 R@5 R@10 R@50
Static 25% 49.26 76.41 84.47 97.03
Dynamic 75% 52.39 79.38 85.49 97.42
Static + Dynamic 100% 53.13 79.93 86.85 97.69
C.6 Effect of modification text length
We analyze the impact of the modification text length, by experimenting with generating multiple candidates per caption pair and selecting the longest modification text, increasing the average length from 23.36 to 33.35 characters. However, as shown in Table A.8, this decreases performance on all three datasets (WebVid-CoVR-Test, CIRR, and FashionIQ). We hypothesize that since our triplets represent a single modification, longer texts tend to be more verbose without improving quality.
Modification text WebVid-CoVR-Test CIRR FashionIQ
avg #chars R@1 MeanR R@1 MeanR R@10 R@50
(ours) 23.36 53.13 79.93 39.06 70.33 27.70 44.63
(longest) 33.35 52.54 78.72 36.67 68.48 24.55 40.49
Table A.8: Increasing the average modification text length in WebVid-CoVR by selecting the longest of multiple generated candidates per caption pair degrades downstream performances on WebVid-CoVR-Test, CIRR, and FashionIQ.
Figure A.8: Training with more target video frames consistently improves performance on WebVid-CoVR-Test.
This shows that while the generated modification texts are not as long on average as the ones from CIRR, our generated texts still provide useful training signal, as evidenced by the state-of-the-art performance of models trained on our dataset when transferred to standard CoIR tasks (CIRR and FashionIQ in Section A.9 and Table A.10). Moreover, the average number of characters of the modification text in WebVid-CoVR and the widely used FashionIQ CoIR benchmark are comparable (see Table 1 of the main paper: 23.36 vs 27.13, respectively).
Mode Pretraining Recall@K $\text{R}_{\text{subset}}$@K
Method Data K=1 K=5 K=10 K=50 K=1 K=2 K=3
Train (CIRR) TIRG (Vo ’19$\dagger$ - 14.61 48.37 64.08 90.03 22.67 44.97 65.14
TIRG+LastConv (Vo ’19) $\dagger$ - 11.04 35.68 51.27 83.29 23.82 45.65 64.55
MAAF (Doods ‘20$\dagger$ - 10.31 33.03 48.30 80.06 21.05 41.81 61.60
MAAF-BERT (Doods ‘20$\dagger$ - 10.12 33.10 48.01 80.57 22.04 42.41 62.14
MAAF-IT (Doods ‘20$\dagger$ - 09.90 32.86 48.83 80.27 21.17 42.04 60.91
MAAF-RP (Doods ‘20$\dagger$ - 10.22 33.32 48.68 81.84 21.41 42.17 61.60
ARTEMIS (Delmas ‘22) - 16.96 46.10 61.31 87.73 39.99 62.20 75.67
CIRPLANT (Liu ‘21$\dagger$ - 19.55 52.55 68.39 92.38 39.20 63.03 79.49
LF-BLIP (Levy 2023) - 20.89 48.07 61.16 83.71 50.22 73.16 86.82
CompoDiff (Gu ‘23) ST18M 22.35 54.36 73.41 91.77 35.84 56.11 76.60
Combiner (Baldrati ‘22) - 33.59 65.35 77.35 95.21 62.39 81.81 92.02
CASE (Levy ‘23) - 48.00 79.11 87.25 97.57 75.88 90.58 96.00
CASE (Levy ‘23) LaSCo 48.68 79.98 88.51 97.49 76.39 90.12 95.86
CASE (Levy ‘23) LaSCo+COCO 49.35 80.02 88.75 97.47 76.48 90.37 95.71
CoVR-BLIP - 49.16 79.76 88.65 97.49 78.12 91.25 96.26
CoVR-BLIP WebVid-CoVR 50.41 80.96 89.35 97.64 77.25 91.04 96.34
[…]
CompoDiff (Gu ‘23) ST18M 19.37 53.81 72.02 90.85 28.96 49.21 67.03
Pic2Word (Saito ‘23) CC3M 23.90 51.70 65.30 87.80 - - -
CASE (Levy ‘23) LaSCo 30.89 60.75 73.88 92.84 60.17 80.17 90.41
CASE (Levy ‘23) LaSCo+COCO 35.40 65.78 78.53 94.63 64.29 82.66 91.61
CoVR-BLIP - 19.95 41.98 52.31 73.54 64.72 83.42 91.95
CoVR-BLIP WebVid-CoVR 39.28 68.22 78.94 94.65 71.23 86.22 94.02
Table A.9: State-of-the-art comparison on the CIRR test set: In Table 3 of the main paper, we presented a subset of the metrics on CIRR dataset due to space constraints. Here, we provide the full set of metrics. Specifically, we add three columns for each $Recall_{\text{subset}}$ metric in CIRR. $\dagger$ denotes results reported by (Liu et al. 2021). For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. CC3M denotes Conceptual Captions 3M (Sharma et al. 2018), ST18M denotes  (Gu et al. 2023), and COCO denotes (Lin et al. 2014).
Mode Pretraining Dress Shirt Toptee Average
Method Data R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
Train (FashionIQ) JVSM (Chen ‘20) - 10.70 25.90 12.00 27.10 13.00 26.90 11.90 26.60
CIRPLANT (Liu ‘21) - 17.45 40.41 17.53 38.81 21.64 45.38 18.87 41.53
TRACE w/BER (Jandial ‘20) - 22.70 44.91 20.80 40.80 24.22 49.80 22.57 46.19
VAL w/GloVe (Chen ‘20) - 22.53 44.00 22.38 44.15 27.53 51.68 24.15 46.61
MAAF (Dodds ‘20) - 23.80 48.60 21.30 44.20 27.90 53.60 24.30 48.80
CurlingNet (Yu ‘20$\ddagger$ - 26.15 53.24 21.45 44.56 30.12 55.23 25.90 51.01
RTIC-GCN (Shin ‘21$\ddagger$ - 29.15 54.04 23.79 47.25 31.61 57.98 28.18 53.09
CoSMo (Lee ‘21) - 25.64 50.30 24.90 49.18 29.21 57.46 26.58 52.31
ARTEMIS (Delmas ‘22) - 27.16 52.40 21.78 43.64 29.20 53.83 26.05 50.29
DCNet (Kim ‘21) - 28.95 56.07 23.95 47.30 30.44 58.29 27.78 53.89
SAC w/BERT (Jandial ‘22) - 26.52 51.01 28.02 51.86 32.70 61.23 29.08 54.70
FashionVLP (Goenka ‘22) - 32.42 60.29 31.89 58.44 38.51 68.79 34.27 62.51
LF-CLIP (Baldrati ‘22) - 31.63 56.67 36.36 58.00 38.19 62.42 35.39 59.03
LF-BLIP (Levy ‘23) - 25.31 44.05 25.39 43.57 26.54 44.48 25.75 43.98
CASE (Levy ‘23) LaSCo 47.44 69.36 48.48 70.23 50.18 72.24 48.79 70.68
CoVR-BLIP - 43.51 67.94 48.28 66.68 51.53 73.60 49.02 70.24
CoVR-BLIP WebVid-CoVR 44.55 69.03 48.43 67.42 52.60 74.31 49.40 70.98
ZS Random - 00.26 01.31 00.16 00.79 00.19 00.95 00.06 00.32
Pic2Word (Saito et al. 2023) CC3M 20.00 40.20 26.20 43.60 27.90 47.40 24.70 43.70
CoVR-BLIP - 13.48 31.96 16.68 30.67 17.84 35.68 16.00 32.77
CoVR-BLIP WebVid-CoVR 21.95 39.05 30.37 46.12 30.78 48.73 27.70 44.63
Table A.10: State-of-the-art comparison on the FashionIQ validation set: In Table 3 of the main paper, we presented a subset of the metrics on FashionIQ dataset due to space constraints. Here, we provide the full set of metrics. Specifically, we add individual categories (Dress, Shirt, and Toptee). For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. $\ddagger$ denotes results reported by (Shin et al. 2021). CC3M denotes Conceptual Captions 3M (Sharma et al. 2018). ZS stands for Zero Shot.
C.7 Optimal number of frames
In Section 4.2 we saw that training with 15 frames further boosts the performance over using only the middle frame. We provide further analysis by incrementally increasing the number of frames for training and testing. We report the average recall on WebVid-CoVR-Test and observe a steady increase in performance with more frames (see Figure A.8).
C.8 Full CoIR metrics
In Table 3 of the main paper, we presented a subset of the metrics on FashionIQ and CIRR datasets due to space constraints. Here, we provide the full set of metrics in Table 3 on CIRR and Table A.9 on FashionIQ. Specifically, we add three columns for each $Recall_{\text{subset}}$ metric in CIRR. For FashionIQ, we add individual categories (Dress, Shirt, and Toptee), as well as the aggregated average results across these categories.
Appendix D Qualitative analysis
In this section, we provide examples of caption filtering (Section D.1), qualitative comparison between different MTG approaches (Section D.2), qualitative examples of our WebVid-CoVR triplets (Section D.3), samples from our manual test set annotation process (Section D.4), qualitative CoVR results on WebVid-CoVR-Test (Section D.5) and CoIR results on CIRR (Section D.6).
D.1 Examples of filtered captions
As described in Section 3.1 of the main paper, we employ a filtering process to select paired captions that facilitate the generation of meaningful training data. In this section, we provide examples of the filtered captions.
Filtering template captions. Upon analyzing the paired captions, we observed that a significant portion of the pairs originated from a small set of template captions. Out of 1.2M distinct caption pairs, approximately 719k ($60\%$) were generated from these template captions. The following examples showcase some of these template captions:
  • Abstract: Abstract color movement tunnel, Abstract color nature background, Abstract color smoke flowing on white background, Abstract colorful paint ink spread explode, Abstract colorful pattern background, Abstract colorful red cement wall background or texture. the camera moves up, Abstract colorful satin background animation, Abstract colorful shiny bokeh background., Abstract colorful smoke on black background, etc
  • Background: Abstract background, Animated backgrounds, Animation, background., Aquarium background, Artistic background, Aurora background, Balloons background, Basketballs background, Beach background, Bluebell background, Bright background, Brush background, Bubbles background, Bubbly background, Celebrate background, Celebratory background, Cg background, Christmas background, Christmas background, Circles background, Color background, Colored background, Colorful background, Colorfull background,, etc.
  • Concept: Brazil high resolution default concept, Brazil high resolution dollars concept, Businessman with advertising hologram concept, Businessman with algorithm hologram concept, Businessman with automation hologram concept, Businessman with bitcoin hologram concept, Businessman with branding hologram concept, Businessman with public relations hologram concept, Close up of an eye focusing on a freelance concept on a futuristic screen., Coins fall into piggy bank painted with flag of ghana. national banking system or savings related conceptual 3d animation, Communication concept, Communication network concept., Communication team concept, Concept of connection, Concept of dancing at disco party. having fun with friends., Concept of education, Concept of geography, Cyber monday concept, etc
  • Flag: Flag of america, Flag of andorra, Flag of aruba, Flag of austria, Flag of azerbaijan, Flag of bahrain, Flag of belarus, Flag of belize, Flag of black, Flag of bolivia, Flag of brazil, Flag of bulgaria, Flag of cameroon, Flag of canada, etc.
Filtering caption pairs with high or low similarity. To ensure the generation of meaningful modifications, we further refine the selection of caption pairs by filtering out those with excessively high or low similarity. Caption pairs with highly similar meanings may result in trivial or unnoticeable modifications. Conversely, pairs with significant dissimilarity can lead to large visual differences that are difficult to describe accurately. We show below some of the filtered captions based on the CLIP text embedding cosine similarity.
  • High similarity: $10\%$ of the pairs have CLIP text similarity above 0.96.
    • Close-up of a tree with green leaves and sunlight
    • Close-up of a tree with green leaves and sunshine
    • Businessman speaking on the phone
    • Businessman talking on the phone
    • Boat on a sea
    • Boat on the sea
  • Low similarity: $2\%$ of the pairs have CLIP text similarity below 0.60.
    • Leaves close-up
    • Peacock, close-up
    • Moon jellyfish
    • Moon night
    • Close up of a lynx
    • Close up of a milkshake
Exclusion of digit differences and out-of-vocabulary words. In order to maintain the high quality and coherence of the generated modification text, we apply additional filtering criteria. Specifically, we exclude caption pairs where the differences between captions are numerical digits (often representing dates) or involve out-of-vocabulary words (using the python libraries wordfreq and enchant) that may hinder the generation process.
  • Difference between the captions is a digit: Approximately $2\%$ of the pairs.
    • 23.09.2015 navigation on the moscow river
    • 07.08.2015 navigation on the moscow river.
    • Light leaks element 190
    • Light leaks element 215
    • Pure silver, shape of granules of pure silver each one is unique 44 (2)
    • Pure silver, shape of granules of pure silver each one is unique 95 (2)
  • Difference in one of the captions has an out-of-vocabulary word: Approximately $7\%$ of the pairs.
    • Businessman writing on hologram desk tech word- bitcoin
    • Businessman writing on hologram desk tech word- crm
    • Mitomycin-c - male doctor with mobile phone opens and touches hologram active ingrident of medicine
    • Oxazepam - male doctor with mobile phone opens and touches hologram active ingrident of medicine
    • Blue forget-me-nots
    • Blue galaxy
Table A.11: Comparison between modification text generation approaches: We provide qualitative examples for a pair of captions, and three methods to generate modification text: (i) rule-based, (ii) prompting-based, (iii) our MTG-LLM finetuning. Rule-based method is limited, for example in the case where the difference text is a preposition (last row), whereas the prompting-based method is prone to hallucinating (e.g., ‘remove iceberg’, ‘change the pose of the runner’). Our approach tends to be the most robust across cases.
Caption1 Happy girl dancing
Caption2 Beautiful girl dancing
Rule-based Make the Happy into Beautiful
Prompting LLM Change girl
MTG-LLM Make her more beautiful
Caption1 Black bird
Caption2 black bear
Rule-based Add bear
Prompting LLM Remove animal
MTG-LLM Turn it into a bear
Caption1 Autumn landscape in the mountains.
Caption2 Winter landscape in the mountains
Rule-based Remove Autumn
Prompting LLM Change scene to winter.
MTG-LLM Change to winter
Caption1 Flying over the ice
Caption2 Flying over the mountains
Rule-based Take out ice and add mountains
Prompting LLM Remove iceberg
MTG-LLM Move the ice to the mountains
Caption1 Dandelion field
Caption2 Rice field
Rule-based Change dandelion for rice
Prompting LLM Change the season
MTG-LLM Turn it into a rice field.
Caption1 Happy woman
Caption2 Running woman
Rule-based Change it to running
Prompting LLM Remove the girl and change the pose of the runner
MTG-LLM Make her a runner
Caption1 Woman working on cellphone at night
Caption2 Woman working on cellphone at outdoor
Rule-based Replace night with outdoor
Prompting LLM Change the woman and add one cellphone
MTG-LLM Make it an outdoor
Caption1 Barber cuts the hair of the client with clipper.
Caption2 Barber cuts the hair of the client with trimmer
Rule-based Take out clipper and add trimmer
Prompting LLM Add or change the hairdresser’s tool
MTG-LLM Change to a trimmer
Caption1 Aerial shot of a lake.
Caption2 Aerial shot above a lake
Rule-based Remove of
Prompting LLM Add a lake to the sequence
MTG-LLM make it an aerial shot of a lake
D.2 Qualitative comparison of MTG approaches
In Section 4.4 of the main paper and Section C.1, we show that finetuning our MTG-LLM works better than a rule-based approach and than few-shot prompting of the LLM. In this section, we provide a qualitative comparison of three different methods for generating modification text: (i) rule-based, (ii) prompting-based, and (iii) our MTG-LLM finetuning. We present examples of paired captions and the corresponding modification texts generated by each method in Table A.11.
Rule-based method. The rule-based method relies on predefined rules to generate modification text. We illustrate an example limitation in the last row of Table A.11, where the difference text is simply a preposition (i.e., ‘of’ vs ‘above’), and the modification text becomes ‘Remove of’. The rule-based method performs well when the modifications follow a specific pattern, but it may struggle with more complex modifications (e.g., ‘tree’ vs ‘trees’ should generate ‘add more trees’ for plurality).
Prompting LLM. The prompting-based method involves using a pretrained language model without finetuning. However, this method is prone to hallucinations and may generate modification text that does not accurately represent the intended difference. For example, in the second example, the prompting LLM suggests removing the term ‘animal’ instead of replacing ‘bird’ with ‘bear’.
MTG-LLM (Our approach). Our MTG-LLM approach utilizes a large language model finetuned on a manually annotated dataset specifically for modification text generation. It tends to be the most robust across different cases.
D.3 Training triplet examples
Figures A.9A.10, and A.11 all show examples of triplets generated using our automatic dataset creation. These examples demonstrate the effectiveness of our approach in generating coherent modification texts for paired videos. This capability serves as a form of data augmentation and increasing the diversity in the training set. In Figure A.12, we show that the dataset is not composed by pairs only, as there are many captions that have many relations between them. Furthermore, in Figure A.13 we show cases where a single caption is associated with multiple videos. This scenario allows us to generate multiple triplets by leveraging the diverse visual content captured in different videos. The triplets shown in the aforemention figures exhibit a wide range of variations, encompassing different themes such as emotions, food, actions, camera edits, gender changes, and time of the day.
D.4 Manual test set annotation
In this section, we further describe the process of manually annotating the test set for our WebVid-CoVR-Test CoVR benchmark, previously discussed in Section 3.2 of the main paper. The annotation process involves presenting the annotator with generated modification texts from three different runs of MTG-LLM, along with three frames each from the query and target videos. The annotator’s task is to evaluate the quality of the modification texts and the suitability of the videos for the CoVR task.
A total of 3.1k triplets were shown for annotation. In Figure A.14 and Figure A.15, we present 10 examples that were considered correct during the annotation, along with the chosen modification texts (marked with a checkmark). These examples demonstrate successful modification texts and appropriate video content for the CoVR task.
On the other hand, in Figure A.15, we show 8 examples that were discarded during the annotation. These examples were rejected either because the modification texts were incorrect or because the videos were deemed unsuitable for the CoVR task due to being either too similar (e.g., bottom left, both videos are showing the same coffe with almost no modification) or too incoherent (e.g., top right example “Make the water a river”).
D.5 Qualitative CoVR results on WebVid-CoVR-Test
In Figure A.16, we show qualitative CoVR results on our manually verified WebVid-CoVR-Test set. We observe that top ranked video frames have high visual and semantic similarity with the queries even when not corresponding to the ground truth (marked with a green border).
D.6 Qualitative CoIR results on the CIRR benchmark
In Figure A.17, we demonstrate qualitative CoIR results of our models trained only on WebVid-CoVR (ZS) and the one further finetuned on CIRR training set (Sup.), tested on the CIRR test set. We observe promising retrieval quality for both models.
Figure A.9: Examples of generated triplets: We illustrate triplet samples (one per row) generated using our automatic dataset creation methodology. Each sample consists of two videos with their corresponding captions (at the bottom of each video) and the generated modification text using our MTG-LLM (in purple).
Figure A.10: Examples of generated triplets (ctd)
Figure A.11: Examples of generated triplets (ctd)
Figure A.12: Generated triplets from multiple similar captions: We can train with as many triplets as pairs of captions with one word difference by generating modification texts using our trained MTG-LLM: she is thinking, Have her look happy, Make the businesswoman pregnant, make her blonde, make her multi-ethnic, Make the woman pregnant, etc.
Figure A.13: Generated triplets with multiple videos: In cases where there are several videos with the same caption, we can generate multiple triplets by leveraging the multiple videos. It can be seen as a way of data augmentation.
Figure A.14: Manual annotation examples (kept): We show samples from WebVid-CoVR-Test which are automatically mined triplets that are marked as correct during the annotation process. Each sample consists of two videos and a set of modification text options (in between each video pair). The chosen modification text is indicated by a checkmark.
Figure A.15: Manual annotation examples (discarded): We show automatically mined triplets that are discarded during the annotation process. Discarded texts include videos that are too similar (bottom left), too dissimilar (bottom right), or have bad modification texts (top left).
Figure A.16: Qualitative CoVR results on WebVid-CoVR-Test: We display the input image and modification text queries on the left, along with the top 3 retrieved videos by our model on the right. Ground-truth is denoted with a green border.
Figure A.17: Qualitative CoIR results on CIRR test set: Given a query image and a modification text, we show our top retrieved videos of our zero-shot (ZS) model trained with WebVid-CoVR and the model finetuned on CIRR ground-truth supervision (Sup.).